Problem 2 Compute the determinant of 1 0 0 B - XId, where Id = 10 1 0 C: 0 0 1 B= 1 0 0 2 1 -2 3 2 1

Also find the Eigenvalues of B
Answer #1

Solution 2 Det (B-n Id), where B- I cd = -14:) (7--1338] B-nId = 1 - 24 2 - 2 1 3 2 Det (B-n Id) = 1-2 Det ( 3 2 I- Dat (B- I